# Tao Yu,  Zongdian Li,  Omar Hashash,  Kei Sakaguchi,  Walid Saad,  and Mérouane Debbah

### v1

@@ Abstract
The concept of digital twin (DT), which enables the creation of a programmable, digital representation of physical systems, is expected to revolutionize future industries and will lie at the heart of the vision of a future smart society, namely, Society 5.0, in which high integration between cyber (digital) and physical spaces is exploited to bring economic and societal advancements. However, the success of such a DT-driven Society 5.0 requires a synergistic convergence of artificial intelligence and networking technologies into an integrated, programmable system that can coordinate networks of DTs to effectively deliver diverse Society 5.0 services. Prior works remain restricted to either qualitative study, simple analysis or software implementations of a single DT, and thus, they cannot provide the highly synergistic integration of digital and physical spaces as required by Society 5.0. In contrast, this paper envisions a novel concept of an Internet of Federated Digital Twins (IoFDT) that holistically integrates heterogeneous and physically separated DTs representing different Society 5.0 services within a single framework and system. For this concept of IoFDT, we first introduce a hierarchical architecture that integrates federated DTs through horizontal and vertical interactions, bridging the cyber and physical spaces to unlock new possibilities. Then, we discuss the challenges of realizing IoFDT, highlighting the intricacies across communication, computing, and AI-native networks while also underscoring potential innovative solutions. Subsequently, we elaborate on the importance of the implementation of a unified IoFDT platform that integrates all technical components and orchestrates their interactions, emphasizing the necessity of practical experimental platforms with a focus on real-world applications in areas like smart mobility.
@@ I Introduction
Digital twins (DTs) are a transformative technology crafting faithful digital representations of physical systems, processes, and dynamics in cyber spaces. This application space enables DTs to span elements ranging within the Internet of Things (IoT) to intelligent transportation systems (ITS), healthcare, and intricate industrial systems across the whole product life-cycle, including design, manufacturing, distribution, and recycling [1]. In essence, a fully realized DT is not just a static blueprint or simulation of a physical system, but a dynamic, high-precision, granular replica of a physical system, including its interactions with the physical system. Thus, DTs symbolize a harmonious blend of physical and digital spaces, which departs from the traditional IoT concept, where interconnectivity is confined among physical objects and data flows mainly in one way from physical to cyber spaces. DTs will be an integral part of future applications ranging from manufacturing to agriculture and enabling the metaverse [2].
@@
DTs will shape the future as a cornerstone of the visionary concept, Society 5.0 [3]. In Society 5.0, the intertwining of cyber and physical spaces becomes a determinant for harmonizing economic advancement with the resolution of societal challenges. In fact, this cyber and physical synergy is facilitated through an intricate web of integration that must consolidate connections between corresponding physical and cyber spaces while performing precise coordination between the DTs of diverse applications. DTs, working in synergy, streamline end-to-end (E2E) operations demanding synchronization across multiple services, such as smart mobility and automated manufacturing. Hence, DTs can establish interconnections and form cooperative clusters in cyber space, collectively modeling and analyzing their corresponding physical systems at varying granularities. Therefore, as a milestone towards enabling Society 5.0, it is necessary to understand the challenges of synergistically interconnecting multiple DTs over wireless networks rather than considering each DT in isolation.
@@
Recent research exploring the synergies between DTs and wireless systems falls into three main categories. The first category (e.g., in [4]), constituting a significant portion of existing research, leverages DTs to create digital representations for wireless systems (e.g., 5G) to enhance network management. While these works are beneficial for wireless systems, they neglect the role of wireless systems in connecting and synchronizing multiple DTs. The second category (e.g., in [5]) studies the deployment of a single, isolated DT within wireless systems, focusing on individual networking and learning schemes. Clearly, these works fall short in extrapolating these network mechanisms for a complex system of interconnected DTs. Finally, some works (e.g., in [6]) focus on virtual representation of twins, neglecting network constraints and primarily tackling software development hurdles. Therefore, these approaches fall short of realizing comprehensive and coordinated E2E DT services in Society 5.0, as current DT technologies lack required features, such as scalability and interconnectivity, that are indispensable in systems of interconnected DTs. Hence, the literature lacks a comprehensive overview that deals with integrating and coordinating multiple DTs across various physical systems with seamless interoperability.
@@
In response to these challenges and to foster interconnections, coordination, and cooperation among DTs while preserving key performance metrics such as scalability, latency, efficiency, and privacy, the main contribution of this paper is to introduce and define a novel, visionary concept called the Internet of Federated Digital Twins (IoFDT). In essence, the IoFDT interconnects a comprehensive and federated network of diverse DTs and smart services, constituting the backbone of Society 5.0. Each DT within an IoFDT is designed for a specific system or service, with federation occurring not only among homogeneous DTs, but broader heterogeneous systems. Moreover, the IoFDT architecture employs dynamic resource allocation and modular services that can adjust system demands, ensuring efficient scalability and performance integrity of the growing network of DTs.
@@
This creates a complex federation of interconnected DTs across various domains, all linked over wireless systems. In this framework, a federated DT is created by leveraging sliced component functions, including sensors, networks, and computing resources integrated within an IoFDT platform. The IoFDT coordinates multiple federated DTs and integrates physically separate, heterogeneous DTs beyond their physical borders. Fig. 1 illustrates an example of the IoFDT vision with several DTs. In this example, the connection between the factory DT and agriculture DT in cyber space enables an automated food processing plant DT. This unified DT in cyber space jointly manages agricultural products from farms and food processing chain in factories. Such integration optimizes the harvest schedule, production plan, quality control, and E2E traceability. Other examples requiring an interconnected IoFDT include various domains: city energy management by connecting power plant DT with user behavior DT; smart supply management by linking logistics DT and manufacturing DT; precision construction by associating construction site DT with construction machinery DT; and smart preventive healthcare by connecting user lifestyle DT to the medical system DT. The IoFDT significantly departs from any existing DT systems that typically rely on isolated, independent DTs. The IoFDT has the potential to revolutionize future societies by providing a backbone for delivering E2E smart services through the seamless integration of digital and physical spaces.
@@
Towards enabling the IoFDT vision, a confluence of AI, networking, computing, and implementation is essential, which introduces several novel challenges that should be addressed. The key contributions of this paper are summarized as follows:
@@
  • We identify multifaceted challenges at the crossroad of communications, artificial intelligence (AI), and computing, and we put forth potential, novel solutions in realizing the IoFDT at enhancing synchronization, cross-layer networking, real-time computing, scalability, security, AI learning, AI-native networks, and generalizability of twins.
  • Transitioning from theoretical concepts to practical realization, we outline the implementation of a unified experimental platform that integrates DT, network, and computing orchestrator to facilitate seamless interactions among diverse DTs. We emphasize real-world testing to validate theoretical models and introduce a comprehensive proof-of-concept (PoC) implementation, with an initial focus on smart mobility, to demonstrate the IoFDT framework’s feasibility and merit.
The rest of this paper is organized as follows. Section II presents the overall IoFDT framework. Section III introduces the enablers and challenges for IoFDT across networking, AI, and computing. Section IV introduces an experimental IoFDT platform with experimental applications. Finally, Section V concludes the paper with future recommendations towards enabling an IoFDT.
@@ II IoFDT: From Individual DTs to Hierarchical, Federated DT Networks
The architecture of the IoFDT consists of a network of interconnected and federated DTs, as illustrated in Fig. 2. This topology overcomes limitations inherent in standalone DT systems, such as lack of interconnectivity, inability to extract multi-layered insights, limited adaptability, and the absence of a collaborative ecosystem.
At the center of our IoFDT lies a series of DTs, each of which serves as a precise digital representation of a physical object or process, bridging cyber and physical spaces. These DTs are strategically arranged in a tiered structure, with each level reflecting intricacies of the DTs’ functionalities and degrees of interaction with their physical counterparts. This hierarchy enables the extraction of rich, multi-layered insights, such as operational details from a an irrigation machine DT, systemic understanding of the irrigation system’s impact on crop growth, and predictive insights on planting cycles and pest control strategies derived from the collective data of farm DTs. DTs, diverse as the systems and services they represent, are not individual but intrinsically linked through a federation. This DT federation is more than a mere technical connection. Instead, it acts as a unifying, logical thread that weaves together distinct DTs, creating a rich tapestry of interactions and information/knowledge exchange. This shared network establishes a collaborative ecosystem within the IoFDT. Participants in this network, representing various DTs, actively exchange real-time data, performance metrics, and AI-derived insights. This continuous exchange not only leverages existing knowledge, but integrates fresh data, refining the system’s collective intelligence and adaptability. The DT federation fosters seamless data synchronization and knowledge sharing among DTs. For instance, a DT representing an irrigation system might adjust water distribution based on insights from another DT monitoring soil moisture. As these DTs interact, they fine-tune their algorithms, ensuring that their digital representations remain aligned with real-world conditions they twin. In the IoFDT framework, interactions are categorized into two main types: horizontal interactions, occurring between DTs of similar hierarchies or functionalities; and vertical interactions, taking place between DTs across different hierarchical levels, bridging granular processes to overarching system dynamics. Next, we describe the horizontal and vertical splits in the IoFDTs:
@@ II-A Horizontal Split
Horizontal interactions within the same levels and clusters of DTs in the hierarchy are fundamental to the IoFDT, because they are the key to improving collaboration, learning, and growth among geographically dispersed DTs. For example, consider the following scenario: A DT provides real-time crop yields and readiness information to a logistics DT, which plans optimal routes for fresh product transportation. Simultaneously, a DT for city-based supermarket chain integrates this information to manage inventory and plan their supply based on fresh product arrival. Meanwhile, the planting or harvest plan on farms will be adjusted according to supermarket demands. This interaction is not only about data exchange, but constitutes continuous mutual learning and growth. Horizontal interaction is strengthened by sophisticated frameworks of real-time data sharing, advanced analytics, and AI-driven insights. It enables DTs to catch and understand other DTs through IoFDT. Furthermore, when multiple DTs, either homogenous or heterogeneous, form functional clusters, the horizontal interactions also facilitate communication across DT clusters. In this interconnected, federated ecosystem, geographical distances and different segments become inconsequential. In a nutshell, the IoFDT effectively shrinks the physical world, by bringing diverse twins closer together in a shared cyber space.
@@ II-B Vertical Split
Having delineated the foundational framework of IoFDT, it is essential to address the underlying technical challenges and enablers over wireless systems. The intricacies and potential solutions at the intersection of communication, AI, and computing will be detailed next.
@@ III Challenges and Enablers for IoFDT: Wireless, Learning, and Computing
Despite benefits of an IoFDT, realizing this vision faces many challenges at the intersection of communication, learning, and computing.
@@ III-A1 Synchronization over the network
Creating and federating multiple DTs requires real-time data accumulation from a multitude of physical objects and processes across vast distances. Meanwhile, modifications in cyber space must be reflected in almost real-time on the physical counterpart. In other words, DTs must be synchronized with the physical space [7]. The twin mandates of IoFDT, i.e., real-time data collection and actuation, present challenges, because they require real-time data transmission between devices of various physical systems and edge/cloud servers through complex, heterogeneous networks. It is also crucial to synchronize interactions between multiple DTs, each DT representing a different aspect or component of a larger system. This further requires maintaining synchronization of contextual data relevance and operational coherence across DTs. Such demands could strain existing communication systems, such as 5G and 6G, while requiring them to constantly adjust to dynamics and requirements of the physical systems, ensuring seamless coordination between various DTs and physical processes. To address synchronization challenges, dynamic network slicing can be developed whereby network resources are adaptively allocated based on each DT’s requirements. This approach, illustrated in Fig. 3, optimizes resource allocation within each slice according to the context and industry-dependent attributes of the physical system. Evolutionary reinforcement learning (ERL)[8], combining reinforcement learning (RL) with evolutionary algorithms, could be employed to evolve slicing policies over time to meet the DT’s dynamic requirements. Also, here one can exploit the notion of semantic communications to push forward less data and facilitate synchronization of DTs [9].
@@ III-A2 Cross-layer networking between DTs
To harness the benefits of DTs and realize IoFDT for Society 5.0, it is necessary to address a series of unique, cross-layer networking challenges inherent in IoFDT. Unlike conventional networks, the IoFDT requires seamless interaction between various networking layers to support the complex functionality of federated DTs. Main challenges include managing asynchronous communications between layers and across different DTs, ensuring data integrity across multiple network protocols, and orchestrating a responsive flow of information that can adapt to the rapidly changing states of physical entities represented by the DTs. These challenges are magnified by diverse natures of the devices involved, each with different communication protocols and data formats, which must be unified within a coherent network structure. The need for real-time analytics and decision-making processes based on contextual information from multiple layers increases the complexity, and requires a more intelligent and flexible network design and management. The tailored solution for these challenges involves network slicing, as shown in Fig. 3, to discern and dynamically accommodate distinct communication and processing requirements at each network layer. This includes the provisioning of network slices with specific cross-layer functionalities like protocol translation, data prioritization, and layer-specific security measures. Moreover, by utilizing a combination of collaborative reinforcement learning (CRL) and transfer learning, the network can continuously learn from data flow patterns and adjust slices in real-time, ensuring optimal cross-layer communication responsive to each DT’s contextual needs. Such an approach enables dynamic and intelligent cross-layer networking that effectively sustains the complex operations of IoFDT to ensure that all network layers work collaboratively for the DT federation.

### v2

@@ Abstract
The concept of digital twin (DT), which enables the creation of a programmable, digital representation of physical systems, is expected to revolutionize future industries and will lie at the heart of the vision of a future smart society, namely, Society 5.0, in which high integration between cyber (digital) and physical spaces is exploited to bring economic and societal advancements. However, the success of such a DT-driven Society 5.0 requires a synergistic convergence of artificial intelligence and networking technologies into an integrated, programmable system that can coordinate DT networks to effectively deliver diverse Society 5.0 services. Prior works remain restricted to either qualitative study, simple analysis or software implementations of a single DT, and thus, they cannot provide the highly synergistic integration of digital and physical spaces as required by Society 5.0. In contrast, this paper envisions a novel concept of an Internet of Federated Digital Twins (IoFDT) that holistically integrates heterogeneous and physically separated DTs representing different Society 5.0 services within a single framework and system. For this concept of IoFDT, we first introduce a hierarchical architecture that integrates federated DTs through horizontal and vertical interactions, bridging cyber and physical spaces to unlock new possibilities. Then, we discuss challenges of realizing IoFDT, highlighting the intricacies across communication, computing, and AI-native networks while also underscoring potential innovative solutions. Subsequently, we elaborate on the importance of the implementation of a unified IoFDT platform that integrates all technical components and orchestrates their interactions, emphasizing the necessity of practical experimental platforms with a focus on real-world applications in areas like smart mobility.
@@ I Introduction
Digital twins (DTs) are a transformative technology crafting faithful digital representations of physical systems, processes, and dynamics in cyber spaces. This application space enables DTs to span elements ranging within the Internet of Things (IoT) to intelligent transportation systems (ITS), healthcare, and intricate industrial systems across the whole product life-cycle, including design, manufacturing, distribution, and recycling [1, 2]. In essence, a fully realized DT is not just a static blueprint or simulation of a physical system, but a dynamic, high-precision, granular replica of a physical system, including its interactions with other physical system. Thus, DTs symbolize a harmonious blend of physical and cyber spaces, which departs from the traditional IoT concept, where interconnectivity is confined among physical objects and data flows mainly in one way from physical to cyber spaces. DTs will be an integral part of future applications ranging from manufacturing to agriculture and enabling the metaverse [3].
@@
DTs will shape the future as a cornerstone of super smart society, dubbed Society 5.0 [4], where the intertwining of cyber and physical spaces drives economic and societal advancement across industries ranging from intelligent transportation to factory automation and robotics. The success of this DT-driven Society 5.0 depends on the seamlessly integration of physical and cyber spaces, ensure precise coordination among DTs and enable them to work together effectively to deliver diverse Society 5.0 services. DTs, working in synergy, streamline end-to-end (E2E) operations demanding synchronization across multiple Society 5.0 services, such as smart mobility and automated manufacturing. Hence, DTs can establish interconnections and form cooperative clusters in cyber space, collectively modeling and analyzing their corresponding physical systems at varying granularities. Therefore, as a milestone towards enabling Society 5.0, it is necessary to understand challenges of synergistically interconnecting multiple DTs over wireless networks rather than considering each DT in isolation.
@@
Recent research exploring synergies between DTs and wireless systems falls into three main categories. The first category (e.g., in [5]), constituting a significant portion of existing research, leverages DTs to create digital representations for wireless systems (e.g., 5G) to enhance network management. While these works are beneficial for wireless systems, they neglect the wireless systems’s role in connecting and synchronizing multiple DTs. The second category (e.g., in [6]) studies the deployment of a single, isolated DT within wireless systems, focusing on individual networking and learning schemes. Clearly, these works fall short in extrapolating these network mechanisms for a complex system of interconnected DTs. Finally, some works (e.g., in [7]) focus on virtual representation of twins, neglecting network constraints and primarily tackling software development hurdles. Therefore, these approaches fall short of realizing comprehensive and coordinated E2E DT services in Society 5.0, as current DT technologies lack required features, such as scalability and interconnectivity, that are indispensable in systems of interconnected DTs. Comprehensive reviews on DTs can be found in [1, 2]. While existing works have explored broader aspects of DTs, they lack a focused exploration on complex integration and coordination of multiple DTs across various physical systems with seamless interoperability, a gap that this paper aims to address.
@@
Building on this identified gap, this paper introduces and defines the Internet of Federated Digital Twins (IoFDT), a novel concept designed to enhance interconnections, coordination, and cooperation among DTs across diverse systems while addressing critical performance metrics such as efficiency, scalability, and latency. In essence, the IoFDT plays a crucial role and constitutes the backbone in realizing Society 5.0 by interconnecting a comprehensive, federated network of diverse DTs and Society 5.0 services. Each DT within an IoFDT in the vision of Society 5.0 is designed for a specific system or service, with federation occurring not only among homogeneous DTs, but in broader heterogeneous systems. Moreover, the IoFDT architecture employs dynamic resource allocation and modular services that can adjust system demands, ensuring efficient scalability and performance integrity of the growing DT network.
@@
This creates a complex federation of interconnected DTs across various domains in Society 5.0, all linked over wireless systems. In this framework, a federated DT leverages sliced component functions, including sensors, networks, and computing resources integrated within an IoFDT platform. The IoFDT coordinates multiple federated DTs and integrates physically separate, heterogeneous DTs beyond their physical borders. Fig. 1 illustrates an example of the IoFDT with several DTs in the vision of Society 5.0. In this example, the connection between the factory DT and agriculture DT in cyber space enables a food processing DT, which manages agricultural products from farms and food processing chain in factories in cyber space. Such integration optimizes harvest schedules, production plans, quality control, and E2E traceability. Other examples requiring an interconnected IoFDT include: city energy management by connecting power plant DT with user behavior DT; smart supply management by linking logistics DT and manufacturing DT; precision construction by associating construction site DT with construction machinery DT; and smart preventive healthcare by connecting user lifestyle DT to medical system DT. The IoFDT significantly departs from conventional DT systems relying on isolated, independent DTs, and could provide a backbone for Society 5.0 to revolutionize future societies and deliver E2E smart services through the seamless integration of cyber and physical spaces.
@@
Towards enabling the IoFDT, a confluence of AI, networking, computing, and implementation is essential, which introduces several novel challenges that should be addressed. This paper’s key contributions are summarized as follows:
@@
  • We identify multifaceted challenges at the crossroad of communications, artificial intelligence (AI), and computing. We put forth potential, novel solutions in realizing the IoFDT at enhancing synchronization, cross-layer networking, real-time computing, scalability, AI learning, AI-native networks, and generalizability of twins.
  • Transitioning from theory to practice, we outline the implementation of a unified experimental platform integrating DT, network, and computing orchestrators to facilitate seamless interactions among diverse DTs. We emphasize real-world testing to validate theoretical models and introduce a comprehensive proof-of-concept (PoC) implementation, initially focused on smart mobility to demonstrate the IoFDT’s feasibility and merit.
The rest of this paper is organized as follows. Section II presents the overall IoFDT framework. Section III introduces enablers and challenges for IoFDT across networking, AI, and computing. Section IV introduces an experimental IoFDT platform with experimental applications. Section V concludes with future recommendations towards enabling IoFDT.
@@ II IoFDT: From Individual DTs to Hierarchical, Federated DT Networks
The IoFDT architecture features a network of interconnected, federated DTs, as illustrated in Fig. 2. This topology overcomes limitations inherent in standalone DT systems, such as lack of interconnectivity, inability to extract multi-layered insights, limited adaptability, and the absence of a collaborative ecosystem.
At the center of IoFDT lies a series of DTs, which serve as precise digital representations of physical objects or processes, bridging cyber and physical spaces. These DTs are strategically arranged in a tiered structure, with each level reflecting intricacies of DTs’ functionalities and degrees of interaction with their physical counterparts. This hierarchy enables the extraction of rich, multi-layered insights, such as operational details from an irrigation machine DT, systemic understanding of the irrigation system’s impact on crop growth, and predictive insights on planting cycles and pest control strategies from collective data of farm DTs. DTs, diverse as systems and services they represent, are not individual but intrinsically linked through a federation. This DT federation is more than a mere technical connection. Instead, it acts as a unifying, logical thread that weaves together distinct DTs, creating a rich tapestry of interactions and information/knowledge exchange. This shared network establishes a collaborative ecosystem within the IoFDT. Participants in this network, representing various DTs, actively exchange real-time data, performance metrics, and AI-derived insights. This continuous exchange not only leverages existing knowledge, but integrates fresh data, refining system’s collective intelligence and adaptability. The DT federation fosters seamless data synchronization and knowledge sharing among DTs. For instance, a DT representing an irrigation system might adjust water distribution based on insights from another DT monitoring soil moisture. As DTs interact, they fine-tune their algorithms, ensuring their digital representations remain aligned with real-world conditions they twin. The IoFDT framework categorizes interactions into two types: horizontal interactions among DTs of similar hierarchies or functionalities; and vertical interactions among DTs across different hierarchical levels, bridging granular processes to overarching system dynamics. Next, we describe the horizontal and vertical splits in the IoFDT:
@@ II-A Horizontal Split
Horizontal interactions within the same DT levels and clusters in the hierarchy are fundamental to the IoFDT for enhancing collaboration, learning, and growth among geographically dispersed DTs. For example, consider the following scenario: A DT provides real-time crop yields and readiness information to a logistics DT, which plans optimal routes for fresh product transportation. Simultaneously, a supermarket chain DT integrates this information to manage inventory and plan their supply based on fresh product arrival. Meanwhile, planting or harvest plans on farms will be adjusted according to supermarket demands. This interaction is not only about data exchange, but constitutes continuous mutual learning and growth. Horizontal interaction is strengthened by sophisticated frameworks of real-time data sharing, advanced analytics, and AI-driven insights. It enables DTs to catch and understand other DTs through the IoFDT. Furthermore, when DTs, either homogeneous or heterogeneous, form functional clusters, the horizontal interactions also facilitate communication across DT clusters. In this interconnected, federated ecosystem, geographical distances and different segments become inconsequential. The IoFDT effectively shrinks the physical world, by bringing diverse twins closer together in a shared cyber space.
@@ II-B Vertical Split
Having delineated the IoFDT’s foundational framework, it is essential to address the underlying technical challenges and enablers over wireless systems. The intricacies and potential solutions at the intersection of communication, AI, and computing will be detailed next.
@@ III Challenges and Enablers for IoFDT: Wireless, Learning, and Computing
Despite IoFDT’s benefits, realizing this vision still faces many challenges.
@@ III-A1 Synchronization over the network
Creating and federating multiple DTs requires real-time data accumulation from a multitude of physical objects and processes across vast distances. Meanwhile, modifications in cyber space must be reflected in almost real-time on the physical counterpart, ensuring DTs must be synchronized with the physical space [8]. The twin mandates of IoFDT, i.e., real-time data collection and actuation, present challenges, because they require real-time data transmission between devices of various physical systems and edge/cloud servers through complex, heterogeneous networks. It is also crucial to synchronize interactions between multiple DTs, each representing different aspects or components of a larger system. This further requires maintaining synchronization of contextual data relevance and operational coherence across DTs. Such demands could strain existing communication systems, such as 5G/6G, while requiring them to constantly adjust to dynamics and requirements of physical systems, ensuring seamless coordination between various DTs and physical processes. To address synchronization challenges, dynamic network slicing can be developed whereby network resources are adaptively allocated based on each DT’s requirements. This approach, illustrated in Fig. 3, optimizes resource allocation within each slice according to the context and industry-dependent attributes of the physical system. Evolutionary reinforcement learning (ERL)[9], combining reinforcement learning (RL) with evolutionary algorithms, could be employed to evolve slicing policies over time to meet DT’s dynamic requirements. Also, here one can exploit semantic communications to push forward less data and facilitate DT synchronization [10].
@@ III-A2 Cross-layer networking between DTs
To harness DTs’ benefits and realize IoFDT for Society 5.0, it is necessary to address a series of unique, cross-layer networking challenges inherent in IoFDT. Unlike conventional networks, IoFDT requires seamless interaction between various networking layers to support complex functionalities of federated DTs. Main challenges include managing asynchronous communications across different layers and DTs, ensuring data integrity across multiple network protocols, and orchestrating responsive information flows that can adapt to the rapidly changing states of physical entities represented by DTs. These challenges are magnified by diverse natures of devices involved, each with different communication protocols and data formats, necessitating a unified coherent network structure. Real-time analytics and decision-making processes based on multi-layer contextual information further increase complexity, requiring more intelligent and flexible network design and management. The tailored solution for these challenges involves network slicing, as shown in Fig. 3, to discern and dynamically accommodate distinct communication and processing requirements at each network layer. This includes the provisioning of network slices with specific cross-layer functionalities like protocol translation, data prioritization, and layer-specific security measures. Moreover, by utilizing a combination of collaborative reinforcement learning (CRL) and transfer learning, the network can continuously learn from data flow patterns and adjust slices in real-time, ensuring optimal cross-layer communication responsive to each DT’s contextual needs. Such an approach enables dynamic and intelligent cross-layer networking that effectively sustains the complex operations of IoFDT to ensure that all network layers work collaboratively for the DT federation.